Shift report 20011221 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Dec 21 2001 - 08:09:23 EST

  • Next message: Flemming Videbaek: "Fw: update for rhic"

    ------------------------------------------------------------
    Date: Fri Dec 21 08:09:23 2001, Shift: 00:00-08:00
    Supervisor: James Norris, with: Flemming Videbaek
    Log book#11 pages 66 - 66
    
    Shift Summary:
    ------------------------------------------------------------
    1. 20:35 - Yellow Queech Link Interlock
       21:45 - Yellow Quench Link Interlock
       22:40 - Preparing to Accelerating 55 bunches
       22:50 - Ramp tuning.  Ramping to injection.
       00:13 - Preparing 55 bunch store for physics
       00:24 - Will ramp for physics store after 
               polarization measurements at injection
       00:47 - Beam has been cogged for physics
    
    		      Triggers 1, 2, 7, 8
    2. PP Data Run 6083 - 759.8K, 308.1K, 6.6K, 10.6K
       PP Data Run 6084 - 233.0K, 144.8K, 3.5K,  5.5K
       PP Data Run 6085 - 177.4K, 120.4K, 3.2K,  5.1K
    
       MRS Angle: 90, FS Angle: 8
    
       Checking timing of trigger counters with beam.
       - The 6 slat behind tofw behaves ok. reasonable
         pedestal, and timing response.
       - The front overlap is made at ~ 17 nsec.
         coinc Left-Right w/o beam ~ 230/sec.
       - Made first attempt for MRS trigger.
         time diff between MRS-Front Left-Right to
         Back 1-6 ~ 114 nsec. Made overlap -
         Fed into trigger 3 (as latch for Min bias run)
         when ever a store succeeds.
    
    3. No Detector Problems.
       Data Run Length Criteria needed - have been 
         stopping at ~5K Trigger 8 because this is
         ~= 100K Trigger 6 from AuAu Data Runs
       Store Dump criteria needed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Dec 21 2001 - 08:09:58 EST